# Prospects of constraining on the polarizations of gravitational waves from binary black holes using space- and ground-based detectors

### v1

@@ VI Conclusion
In this paper, we investigate the expected targets for constraining GW polarization using space- and ground-based detectors within the ppE framework. Specifically, we adopt a model-independent ppE framework that incorporates a GW waveform with all six polarization modes, adhering to current GR test results. For space-based detectors, we consider LISA, Taiji, and TianQin, along with alternative orbital configurations, both individually and in network scenarios. For ground-based detectors, we include the currently operational LVK and the third-generation detector ET, evaluated both individually and in network. Our analysis focuses on detector performance across different polarization modes, emphasizing the response function’s perspective. Furthermore, by simulating three typical masses of MBHB and SBBH, we use the FIM to quantify ppE parameter constraints for both space- and ground-based detectors, presenting the results under different combinations. Furthermore, we explore how multiband and multimessenger observations enhance ppE constraints, offering a comprehensive analysis from diverse angles.
@@
For space-based detectors, Taiji provides the best constraints on GW polarization modes, with LISA performing better than TianQin. Specifically, TianQin’s constraints on scalar modes are significantly weaker than those of LISA and Taiji, while the uncertainties in other ppE parameters among the three detectors differ by less than an order of magnitude. In network scenarios, the combination of LISA+TJm performs slightly better than other combinations, and the network of all three detectors outperforms two-detector networks.
@@
For ground-based detectors, the differences in constraints on frequency-related ppE parameters are minor, but LIGO’s results for vector modes are substantially better than those of Virgo and KAGRA, approaching the level of ET. In network scenarios, the LVK combination does not surpass the results of ET alone, with the LVK+ET network providing the best constraints. Overall, ground-based detectors are unable to distinguish between breathing and longitudinal modes and are significantly less capable of constraining scalar modes compared to space detectors. Other than this, the constraints on ppE parameters from space- and ground-based detectors are comparable.
@@
For multiband observation, it can effectively break the degeneracies of ground-based detectors in breathing and longitudinal modes, addressing the inability of ground detectors to distinguish scalar modes. They also improve the ability of Virgo and KAGRA to constrain vector modes, reaching the level of LIGO, and offer certain improvements for other ppE parameters. Moreover, TianQin brings the greatest improvement, followed by Taiji, and LISA has the least improvement.
@@
For multimessenger observation, it provides the most significant enhancement for TianQin, bringing its results for vector modes close to those of LISA. For scalar modes, LISA and Taiji see greater improvements. ET shows a larger improvement compared to LIGO, and Virgo and KAGRA experience substantial enhancements, with their results for vector modes approaching the level of LIGO. Overall, both multiband and multimessenger observations can enhance GW detection results to a certain extent, leading to better constraints on GW polarization modes.
@@
In future research, we plan to approach the study from several aspects. First, we will use longer-duration GW signals, which can contain more information, and consider the non-light-speed propagation caused by different modes, where vector and scalar modes may arrive before or simultaneously with tensor modes [70]. Second, we aim to delve into the effects brought by Time-Delay Interferometry (TDI) technology, as different TDI combinations may result in subtle differences in the outcomes [71]. Additionally, we will consider various sources, examining the impact of ppE parameters on unequal mass BBH systems. Lastly, in the context of multimessenger observations, we will study the performance of EM detectors under realistic conditions, providing results that are more aligned with actual scenarios to enhance GW observations. In summary, through these in-depth studies, we aim to advance the research on GW polarization modes, evaluate detector performance more comprehensively and in greater detail, and provide more valuable results.
@@ Acknowledgements.
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the National Natural Science Foundation of China (Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103).

### v2

@@ VI Conclusion
In this paper, we investigate the expected targets for constraining GW polarization using space- and ground-based detectors within the ppE framework. Specifically, we adopt a model-independent ppE framework that incorporates a GW waveform with all six polarization modes, adhering to current GR test results. The space-based detectors considered in this study include LISA, Taiji, and TianQin, along with alternative orbital configurations for individual use or in network scenarios. As for the ground-based detectors, we examine the currently operational LVK and the third-generation detector ET, evaluating them both individually and in network configurations. Our analysis focuses on the performance of detectors across different polarization modes, with an emphasis on the perspective of the response function. Additionally, we employ the FIM to quantify ppE parameter constraints for both space- and ground-based detectors by simulating three typical masses of MBHB and SBBH. We present the results under various combinations and further explore how multiband and multimessenger observations enhance ppE constraints, providing a comprehensive analysis from diverse perspectives.
@@
For space-based detectors, Taiji provides the most stringent constraints on GW polarization modes, with LISA outperforming TianQin. Specifically, TianQin exhibits significantly weaker constraints on scalar modes compared with LISA and Taiji, but the uncertainties in other ppE parameters are similar between the two detectors. In network scenarios, the combination of LISA+TJm slightly surpasses other combinations in performance, and a network comprising all three detectors outperforms networks consisting of only two detectors.
@@
For ground-based detectors, the differences in constraints on frequency-related ppE parameters are negligible, while LIGO’s results for vector modes outperform those of Virgo and KAGRA, approaching the performance level of ET. In network scenarios, the LVK combination does not surpass the results of ET alone, with the LVK+ET network providing the most optimal constraints.. Overall, ground-based detectors lack discrimination between breathing and longitudinal modes and exhibit significantly lower capability in constraining scalar modes compared with space detectors. Apart from this distinction, the constraints on ppE parameters from both space- and ground-based detectors are comparable.
@@
For multiband observation, it can effectively resolves the degeneracies encountered by ground-based detectors in breathing and longitudinal modes, addressing their inability to discern scalar modes. This approach also enhances the capability of Virgo and KAGRA to constrain vector modes, reaching a level comparable to LIGO and offering certain improvements for other ppE parameters. Furthermore, TianQin exhibits the most significant enhancement, followed by Taiji, with LISA demonstrating comparatively lesser improvement.
@@
For multimessenger observation, it provides the most significant enhancement for TianQin, bringing its results for vector modes close to those of LISA. Regarding scalar modes, both LISA and Taiji witness greater improvements. ET demonstrates a larger improvement compared with LIGO, while Virgo and KAGRA experience substantial enhancements with their results for vector modes approaching the level of LIGO. Overall, both multiband and multimessenger observations can augment GW detection outcomes to a certain extent, leading to improved constraints on GW polarization modes.
@@
The future research will be approached from multiple perspectives. Firstly, we will utilize longer-duration GW signals, which possess a greater capacity for information content. Additionally, we will consider the propagation of GWs at speeds other than that of light due to different modes. It is worth noting that vector and scalar modes may arrive before or simultaneously with tensor modes [77]. Secondly, our objective is to delve into the effects brought about by time-delay interferometry (TDI) technology. This entails exploring the subtle differences in outcomes resulting from various TDI combinations [78]. Moreover, we will examine diverse sources and assess the impact of ppE parameters on BBH systems with unequal masses. Thirdly, our plan involves employing specific modified theories of gravity in conjunction with existing GW events to yield more stringent constraint results [79, 80, 81, 82, 83, 84]. Lastly, within the context of multimessenger observations, we aim to investigate the performance of EM detectors under realistic conditions. This approach allows us to provide results that align more closely with actual scenarios and enhance GW observations. The aim of these in-depth studies is to advance the research on GW polarization modes, comprehensively and meticulously evaluate detector performance, and provide more valuable results.
@@ Acknowledgements.
This work was supported by the National Key Research and Development Program of China (Grant No. 2023YFC2206702), the National Natural Science Foundation of China (Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103).